# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v4

@@ 8. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. To tackle this, we propose a novel solution—an LLM-powered experts-in-the-loop chatbot framework, that utilizes retrieval-augmented generation over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
@@ Acknowledgements.
- Beisecker and Beisecker (1990) Analee E. Beisecker and Thomas D. Beisecker. 1990. Patient Information-Seeking Behaviors When Communicating with Doctors. Medical Care 28, 1 (1990), 19–28. http://www.jstor.org/stable/3765618
@@
- Bérubé et al. (2021) Caterina Bérubé, Zsolt Ferenc Kovacs, Elgar Fleisch, and Tobias Kowatsch. 2021. Reliability of Commercial Voice Assistants’ Responses to Health-Related Questions in Noncommunicable Disease Management: Factorial Experiment Assessing Response Rate and Source of Information. Journal of Medical Internet Research 23, 12 (2021), e32161.
@@
- Desai and Chin (2023) Smit Desai and Jessie Chin. 2023. OK Google, Let’s Learn: Using Voice User Interfaces for Informal Self-Regulated Learning of Health Topics among Younger and Older Adults. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 847, 21 pages. https://doi.org/10.1145/3544548.3581507
@@
- Findlater et al. (2009) Leah Findlater, Ravin Balakrishnan, and Kentaro Toyama. 2009. Comparing Semiliterate and Illiterate Users’ Ability to Transition from Audio+text to Text-Only Interaction. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Boston, MA, USA) (CHI ’09). Association for Computing Machinery, New York, NY, USA, 1751–1760. https://doi.org/10.1145/1518701.1518971
@@
- Fogel and Kvedar (2018) Alexander L Fogel and Joseph C Kvedar. 2018. Artificial intelligence powers digital medicine. NPJ digital medicine 1, 1 (2018), 5.
@@
- Henry et al. (2022) Katharine E Henry, Rachel Kornfield, Anirudh Sridharan, Robert C Linton, Catherine Groh, Tony Wang, Albert Wu, Bilge Mutlu, and Suchi Saria. 2022. Human–machine teaming is key to AI adoption: clinicians’ experiences with a deployed machine learning system. NPJ digital medicine 5, 1 (2022), 97.
@@
- Ma et al. (2023) Xiumei Ma, Yunxing Liu, Pengfei Zhang, Rongtao Qi, and Fanbo Meng. 2023. Understanding online health information seeking behavior of older adults: A social cognitive perspective. Frontiers in Public Health 11 (2023), 1147789.
@@
- Roller et al. (2020) Stephen Roller, Emily Dinan, Naman Goyal, Da Ju, Mary Williamson, Yinhan Liu, Jing Xu, Myle Ott, Kurt Shuster, Eric M Smith, et al. 2020. Recipes for building an open-domain chatbot. arXiv preprint arXiv:2004.13637 (2020).
@@
- Williams and Moser (2019) Michael Williams and Tami Moser. 2019. The art of coding and thematic exploration in qualitative research. International Management Review 15, 1 (2019), 45–55.
- Wutz et al. (2023) Maximilian Wutz, Marius Hermes, Vera Winter, and Juliane Köberlein-Neu. 2023. Factors influencing the acceptability, acceptance, and adoption of conversational agents in health care: integrative review. Journal of Medical Internet Research 25 (2023), e46548.
@@ A.1.1. Cataract Bot LLM Prompts
A.2. Chatbot Usability Form
CataractBot understands me well.
CataractBot’s responses are easy to understand.
CataractBot responds quickly.
CataractBot seems to have a good grasp of medical knowledge.
CataractBot is kind and helpful.
CataractBot is easy to use.
A.3. Participant Demography
@@
| Id | Role | Age | Sex | Languages | Highest Education | Experience |
| D1 | Operating Doctor | 42 | F | English, Hindi, Kannada, Tamil, Telugu | Masters | 10+ years |
| D2 | Operating Doctor | 44 | F | English, Hindi, Kannada, Tamil, Telugu, Malyalam | Masters | 15+ years |
| D3 | Operating Doctor | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| D4 | Escalation Doctor; KB update expert | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| C1 | Coordinator | 33 | F | English, Hindi, Kannada, Tamil, Telugu | Bachelors | 10+ years |
| C2 | Escalation Coordinator | 36 | F | English, Hindi, Kannada, Tamil, Malyalam | Bachelors | 15+ years |

### v5

@@ 8. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. We propose a novel solution—an LLM-powered experts-in-the-loop chatbot framework, that utilizes retrieval-augmented generation over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 information seekers (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
@@ Acknowledgements.
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (2006), 77–101. https://doi.org/10.1191/1478088706qp063oa arXiv:https://www.tandfonline.com/doi/pdf/10.1191/1478088706qp063oa
@@
- eMed (2024) eMed. 2024. Babylon Health. Retrieved Sept 1, 2024 from https://www.emed.com/uk
@@
- Kreuter and McClure (2004) Matthew W Kreuter and Stephanie M McClure. 2004. The role of culture in health communication. Annu. Rev. Public Health 25, 1 (2004), 439–455.
@@
- Li et al. (2025) David Li, Kartik Gupta, Mousumi Bhaduri, Paul Sathiadoss, Sahir Bhatnagar, and Jaron Chong. 2025. Comparative diagnostic accuracy of GPT-4o and LLaMA 3-70b: Proprietary vs. open-source large language models in radiology. Clinical Imaging 118 (2025), 110382. https://doi.org/10.1016/j.clinimag.2024.110382
@@
- McGhee et al. (2020) Charles NJ McGhee, Jie Zhang, and Dipika V Patel. 2020. A perspective of contemporary cataract surgery: the most common surgical procedure in the world. Journal of the Royal Society of New Zealand 50, 2 (2020), 245–262.
@@
- Sun et al. (2013) Si Sun, Xiaomu Zhou, Joshua C. Denny, Trent S. Rosenbloom, and Hua Xu. 2013. Messaging to your doctors: understanding patient-provider communications via a portal system. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Paris, France) (CHI ’13). Association for Computing Machinery, New York, NY, USA, 1739–1748. https://doi.org/10.1145/2470654.2466230
@@
- Thirunavukarasu et al. (2023) Arun James Thirunavukarasu, Darren Shu Jeng Ting, Kabilan Elangovan, Laura Gutierrez, Ting Fang Tan, and Daniel Shu Wei Ting. 2023. Large language models in medicine. Nature medicine 29, 8 (2023), 1930–1940.
@@ A.1.1. CataractBot LLM Prompts
A.1.2. CataractBot’s Support for Multimodal Communication
See Figure 4.
A.2. Participant Demography
@@
| Id | Role(s) | Age | Sex |
Experience |
| D1 | Operating doctor | 42 | F | Masters | 10+ years | 10-20 | No | Yes | ||||||||
| D2 | Operating doctor | 44 | F | Masters | 15+ years | 10-20 | No | Yes | ||||||||
| D3 | Operating doctor | 46 | M | Masters | 20+ years | 30-40 | Yes | Yes | ||||||||
| D4 | Escalation doctor, Knowledge base expert | 46 | M | Masters | 20+ years | 15-20 | Yes | Yes | ||||||||
| D5 | Doctor | 30 | M | Masters | 6+ years | 10 | Yes | No | ||||||||
| D6 | Doctor | 29 | F | Masters | 2+ years | 1-2 | Yes | No | ||||||||
| C1 | Operating patient coordinator | 33 | F | Bachelors | 10+ years | N/A | Yes | Yes | ||||||||
| C2 | Escalation patient coordinator | 36 | F | Bachelors | 15+ years | N/A | No | Yes | ||||||||
| C3 | Patient coordinator | 45 | F | Bachelors | 10+ years | N/A | Yes | No |
A.3. Chatbot Usability Evaluation Form
CataractBot understands me well.
CataractBot’s responses are easy to understand.
CataractBot responds quickly.
CataractBot seems to have a good grasp of medical knowledge.
CataractBot is kind and helpful.
CataractBot is easy to use.
A.4. Summary of Key Findings
| Research Q | Theme | Finding (Information Seekers) | Finding (Experts) | ||||||||
| RQ1: Information Needs | Reasons for usage |
• Commitment to patients | ||||||||
| RQ2: Features | LLM-powered |
| Experts-in-the-loop |
• Hard to decipher English transcriptions | |||||||||
| Knowledge base |
| RQ3: Workflow | Workflows |
See Table 5